# Tiled Flash Linear Attention: More Efficient Linear RNN and xLSTM Kernels

### v2

@@ A.2 Other Related Work
Optimized, hardware-aware implementations enable the exploration of new primitives or new model architectures. FlashRNN (Pöppel et al., 2025) introduces a framework of IO-aware optimized CUDA kernels in order to simplify research on traditional, non-parallelizable RNNs. Mamba (Gu & Dao, 2024) enables large scale language modeling experiments (Waleffe et al., 2024) with an efficient parallel scan algorithm in their optimized CUDA kernels. FlashFFTConv (Fu et al., 2024) provides efficient implementations for FFT convolutions for modern hardware by reducing IO and leveraging specialized matrix multiply units. DeltaNet (Yang et al., 2024c, a) introduces an efficient algorithm for training linear Transformers with the delta rule (DeltaNet) (Schlag et al., 2021), which enables to scale up DeltaNet to standard language modeling settings.
Our TFLA kernel algorithm provides an effective method to balance the runtime and memory for linear RNN kernels based on their chunkwise-parallel formulation, paving the way to even larger model training setups.
@@
A.3 Application of TFLA to other Linear RNNs
@@
By defining the new values $\bm{V}^{*(k)}=\bm{U}^{(k)}-\bm{W}^{(k)}\bm{C}_{k-1}$ we can recover the core formula of TFLA for the intra-chunk parallel forward pass (see Equation (20)) to which we can apply our TFLA tiling strategy from Section 3.2. For computing the matrices $\bm{U}^{(k)}$, $\bm{W}^{(k)}$ and the new values $\bm{V}^{*(k)}$ one could follow the same strategy as in DeltaNet, where $\bm{U}^{(k)}$ and $\bm{W}^{(k)}$ are computed in a separate kernel before the recurrent kernel. Then, the new values $\bm{V}^{*(k)}$ are computed and stored in HBM in the recurrent kernel together with the memory states $\bm{C}_{k}$.
@@ 10 L = chunk_size
$\widetilde{\bm{\mathrm{D}}}^{(k)}$ =$=\begin{cases}-\infty&\text{for}\;i<j\\ \bm{\mathrm{b}}_{k}-\bm{\mathrm{b}}_{k}^{\top}+{\overline{\bm{\mathrm{i}}}^{(k)}}^{\top}&\text{for}\;i\geqslant j\end{cases}$ (46)
@@ B.3 Chunkwise-Parallel Backward Pass
[…]
@@
[…]
@@
where we use the logarithmic forget gates $\bar{\bm{\mathrm{f}}}=\log\sigma(\tilde{\bm{\mathrm{f}}})$. We denote the dimensions as $L_{q}$ and $L_{kv}$ for the query and key-value dimensions, respectively. Omitting the masking operation, we compute the gate matrix as
@@
[…]
@@ Appendix C Extended Tiled Flash Linear Attention
C.1 Tiled Computation
@@
C.2 TFLA Forward Pass
@@
Note that the loop in line 27 of Algorithm 1 is the same as the loop in line 6. In both loops we load the same blocks of the matrix $\bm{Q}$. Fusing these loops would avoid loading this data twice. Unfortunately, fusing these loops efficiently is problematic due to the online computation of the maximum $\bm{m}_{old}$ and $\bm{m}_{new}$ in the loop in line 4 and the dependence of $\bm{m}^{(\text{combine})}_{k}$ and $\overline{\bm{\mathrm{b}}}_{k}$ on the final $\bm{m}_{new}$ (see Appendix D.1 and C.1).
@@
0:  Matrices $\bm{Q}\in\mathbb{R}^{L_{hq}\times d_{qk}}$, $\bm{K}\in\mathbb{R}^{L_{kv}\times d_{qk}}$, $\bm{V}\in\mathbb{R}^{L_{kv}\times d_{hv}}$.
@@
1:  Initialize $\bm{m}_{old},\,\bm{m}_{new}\in\mathbb{R}^{L_{q}}$ to $\infty$- ∞ in SRAM.
@@
2:  Initialize accumulators $\bm{\mathrm{H}}_{\text{intra}}\in\mathbb{R}^{B_{Lhq}\times B_{dv}}$ and $\bm{\mathrm{n}}^{(\text{intra})}\in\mathbb{R}^{B_{Lhq}}$ in SRAM.
3:  Load $\bm{\mathrm{b}}_{k}^{(q)}\in\mathbb{R}^{B_{Lhq}}$ from HBM to SRAM.
4:  for $i=1$ to $\left\lfloor\frac{(i_{Lq}+1)\cdot B_{Lhq}}{B_{Lkv}}\right\rfloor$ do
5:     Initialize accumulator $\bm{\mathrm{S}}^{(i)}\in\mathbb{R}^{B_{Lhq}\times B_{Lkv}}$ in SRAM.
6:     for $j=1$ to $\left\lceil\frac{d_{qk}}{B_{dqk}}\right\rceil$ do
7:        Load $\bm{Q}^{(j)}\in\mathbb{R}^{B_{Lhq}\times B_{dqk}}$ and $\bm{K}^{(j)}\in\mathbb{R}^{B_{Lkv}\times B_{dqk}}$ from HBM to SRAM.
8:        Accumulate $\bm{\mathrm{S}}^{(i)}\mathrel{+}=\bm{Q}^{(j)}{\bm{K}^{(j)}}^{\top}$.
9:     end for
10:     Load $\bm{\mathrm{b}}_{k}^{(kv)}\in\mathbb{R}^{B_{Lkv}}$ and $\bm{\mathrm{i}}_{k}^{(kv)}\in\mathbb{R}^{B_{Lkv}}$ from HBM to SRAM.
11:     Compute $\widetilde{\bm{\mathrm{D}}}^{(i)}=\bm{\mathrm{b}}_{k}^{(q)}-{\bm{\mathrm{b}}_{k}^{(kv)}}^{\top}+{\bm{\mathrm{i}}_{k}^{(kv)}}^{\top}\in\mathbb{R}^{B_{Lhq}\times B_{Lkv}}$.
12:     if $i\cdot B_{Lkv}\geqslant i_{Lq}\cdot B_{Lhq}$ then
13:        Apply causal mask to $\widetilde{\bm{\mathrm{D}}}^{(i)}$.
14:     end if
15:     Compute $\bm{m}_{new}=\mathrm{maximum}\{\bm{m}_{old},\,\mathrm{rowmax}\;\widetilde{\bm{\mathrm{D}}}^{(i)}\}$.
16:     Compute $\bm{\mathrm{D}}^{(i)}=\exp(\widetilde{\bm{\mathrm{D}}}^{(i)}-\bm{m}_{new})$.
17:     Compute $\overline{\bm{\mathrm{S}}}^{(i)}=\frac{1}{\sqrt{d_{qk}}}\bm{\mathrm{S}}\odot\bm{\mathrm{D}}^{(i)}$.
18:     Load $\bm{V}^{(i)}\in\mathbb{R}^{B_{Lkv}\times B_{dhv}}$ for Block $i_{dhv}$ from HBM to SRAM.
19:     Accumulate $\bm{\mathrm{H}}_{\text{intra}}=\exp(\bm{m}_{old}-\bm{m}_{new})\cdot\bm{\mathrm{H}}_{\text{intra}}+\overline{\bm{\mathrm{S}}}^{(i)}\,\bm{V}$.
20:     Accumulate $\bm{\mathrm{n}}^{(\text{intra})}=\exp(\bm{m}_{old}-\bm{m}_{new})\cdot\bm{\mathrm{n}}^{(\text{intra})}+\overline{\bm{\mathrm{S}}}^{(i)}\,\bm{1}$.
21:     Update $\bm{m}_{old}=\bm{m}_{new}$.
22:  end for
@@
23:  Load ${m}^{(\text{inter})}_{k-1}\in\mathbb{R}$ from HBM to SRAM.
24:  Compute $\bm{m}^{(\text{combine})}_{k}=\mathrm{maximum}\left\{\bm{\mathrm{b}}_{k}^{(q)}+{m}^{(\text{inter})}_{k-1},\,\bm{m}_{new}\right\}$.
25:  Compute $\overline{\bm{\mathrm{b}}}_{k}=\exp\left(\bm{\mathrm{b}}_{k}^{(q)}+{m}^{(\text{inter})}_{k-1}-\bm{m}^{(\text{combine})}_{k}\right)$.
26:  Initialize accumulators $\bm{\mathrm{H}}_{\text{inter}}\in\mathbb{R}^{B_{Lhq}\times B_{dhv}}$ for Block $i_{dhv}$ and $\bm{\mathrm{n}}^{(\text{inter})}\in\mathbb{R}^{B_{Lhq}}$ in SRAM.
@@
27:  for $j=1$ to $\left\lceil\frac{d_{qk}}{B_{dqk}}\right\rceil$ do
28:     Load $\bm{Q}^{(j)}\in\mathbb{R}^{B_{Lhq}\times B_{dqk}}$ and $\bm{C}_{k-1}^{(j)}\in\mathbb{R}^{B_{dqk}\times B_{dhv}}$ for Block $i_{dhv}$ from HBM to SRAM.
29:     Compute $\overline{\bm{Q}}^{(j)}=\frac{1}{\sqrt{d_{qk}}}\,\bm{Q}^{(j)}\odot\bm{\mathrm{b}}_{k}^{(q)}$.
30:     Accumulate $\bm{\mathrm{H}}_{\text{inter}}\mathrel{+}=\overline{\bm{Q}}^{(j)}\,\bm{C}_{k-1}^{(j)}$.
31:     Load $\bm{\bm{n}}_{k-1}^{(j)}\in\mathbb{R}^{B_{dqk}}$.
32:     Accumulate $\bm{\mathrm{n}}^{(\text{inter})}\mathrel{+}=\overline{\bm{Q}}^{(j)}\,\bm{\bm{n}}_{k-1}^{(j)}$.
33:  end $\triangleright$ Combine inter- and intra-chunk contributions
34:  Compute $\bm{\mathrm{H}}^{(\text{comb})}=\bm{\mathrm{H}}_{\text{intra}}+\exp(\bm{m}_{new}-\bm{m}^{(\text{combine})}_{k})\,\bm{\mathrm{H}}_{\text{inter}}$.
35:  Compute $\bm{\mathrm{n}}^{(\text{comb})}=\bm{\mathrm{n}}^{(\text{intra})}+\exp(\bm{m}_{new}-\bm{m}^{(\text{combine})}_{k})\,\bm{\mathrm{n}}^{(\text{inter})}$.
36:  Compute $\bm{\mathrm{H}}^{(k)}=\frac{\bm{\mathrm{H}}^{(\text{comb})}}{\max\left\{|\bm{\mathrm{n}}^{(\text{comb})}|,\exp(-\bm{m}^{(\text{combine})}_{k})\right\}}$.
37:  Store $\bm{\mathrm{H}}^{(k)}$, $\bm{\mathrm{n}}^{(\text{comb})}$ and $\bm{m}^{(\text{combine})}_{k}$ to HBM.
C.3 TFLA Backward Pass
@@ D.1 Stabilization of the Exponential Input Gate
[…]
@@
$\widetilde{\bm{C}}_{t}$ $=\bm{C}_{t}\,\exp(-m_{t})$ $\rightarrow\ \widetilde{\bm{C}}_{t-1}=\bm{C}_{t-1}\,\exp(-m_{t-1})$ $\Leftrightarrow\ {\bm{C}}_{t-1}=\widetilde{\bm{C}}_{t-1}\,\exp(m_{t-1})$ (110)
$\widetilde{\bm{n}}_{t}$ $=\bm{n}_{t}\,\exp(-m_{t})$ $\rightarrow\ \widetilde{\bm{n}}_{t-1}\ =\bm{n}_{t-1}\,\exp(-m_{t-1})$ $\Leftrightarrow\ {\bm{n}}_{t-1}\ =\widetilde{\bm{n}}_{t-1}\,\exp(m_{t-1})$

### v3

@@ A.1 Other Related Work
Optimized, hardware-aware implementations enable the exploration of new primitives or new model architectures. FlashRNN Pöppel et al. (2025) introduces a framework of IO-aware optimized CUDA kernels in order to simplify research on traditional, non-parallelizable RNNs. Mamba (Gu & Dao, 2024) enables large scale language modeling experiments (Waleffe et al., 2024) with an efficient parallel scan algorithm in their optimized CUDA kernels. FlashFFTConv (Fu et al., 2024) provides efficient implementations for FFT convolutions for modern hardware by reducing IO and leveraging specialized matrix multiply units. DeltaNet Yang et al. (2024c, a) introduces an efficient algorithm for training linear Transformers with the delta rule (DeltaNet) (Schlag et al., 2021), which enables to scale up DeltaNet to standard language modeling settings.
Our TFLA kernel algorithm provides an effective method to balance the runtime and memory for linear RNN kernels based on their chunkwise-parallel formulation, paving the way to even larger model training setups, e.g. for multi-billion parameter xLSTM models (Beck et al., 2025).
@@
A.2 Application of TFLA to other Linear RNNs
@@
By defining the new values $\bm{V}^{*(k)}=\bm{U}^{(k)}-\bm{W}^{(k)}\bm{C}_{k-1}$ we can recover the core formula of TFLA for the intra-chunk parallel forward pass (see Equation (12)) to which we can apply our TFLA tiling strategy from Section 3. For computing the matrices $\bm{U}^{(k)}$, $\bm{W}^{(k)}$ and the new values $\bm{V}^{*(k)}$ one could follow the same strategy as in DeltaNet, where $\bm{U}^{(k)}$ and $\bm{W}^{(k)}$ are computed in a separate kernel before the recurrent kernel. Then, the new values $\bm{V}^{*(k)}$ are computed and stored in HBM in the recurrent kernel together with the memory states $\bm{C}_{k}$.
@@ 10 L = chunk_size
$\widetilde{\bm{\mathrm{D}}}^{(k)}$ $=\begin{cases}-\infty&\text{for}\;i<j\\ \bm{\mathrm{b}}_{k}-\bm{\mathrm{b}}_{k}^{\top}+{\overline{\bm{\mathrm{i}}}^{(k)}}^{\top}&\text{for}\;i\geqslant j\end{cases}$ (36)
@@ B.3 Chunkwise-Parallel Backward Pass
$\delta\bm{K}^{(k)}_{\text{inter}}$ $=\delta\overline{\bm{K}}^{(k)}\odot\overline{\bm{a}}_{k}\ \bm{1}^{\top}$ (52)
@@
$\delta\overline{\bm{b}}_{k}$ $=(\delta\overline{\bm{Q}}^{(k)}\odot\frac{\bm{Q}^{(k)}}{\sqrt{d_{qk}}})\ \bm{1}$ (61)
@@
$\delta\bm{b}_{k}^{(q)}$ $=\delta\bm{\mathrm{D}}^{(k)}\ \bm{1}_{kv}$ (75)
$\delta\bm{b}_{k}^{(kv)}$ $=-{\delta\bm{\mathrm{D}}^{(k)}}^{\top}\,\bm{1}_{q}$ (76)
$\delta\overline{\bm{\mathrm{i}}}_{kv}^{(k)}$ $={\delta\bm{\mathrm{D}}^{(k)}}^{\top}\,\bm{1}_{q}.$ (77)
@@ Appendix C Extended Tiled Flash Linear Attention
C.1 GPU Fundamentals
We review the GPU fundamentals for writing efficient kernels. Since we perform our experiments on NVIDIA GPUs, our review is targeted towards NVIDIA’s terminology, though the principles also apply to other hardware. For a more extensive overview we refer to (Spector et al., 2024).
GPU Overview.
C.2 Tiled Computation
@@
C.3 TFLA Forward Pass
@@
Note that the loop in line 27 of Algorithm 1 is the same as the loop in line 6. In both loops we load the same blocks of the matrix $\bm{Q}$. Fusing these loops would avoid loading this data twice. Unfortunately, fusing these loops efficiently is problematic due to the online computation of the maximum $\bm{m}_{old}$ and $\bm{m}_{new}$ in the loop in line 4 and the dependence of $\bm{m}^{(\text{combine})}_{k}$ and $\overline{\bm{\mathrm{b}}}_{k}$ on the final $\bm{m}_{new}$ (see Appendix D.1 and C.2).
@@
1:Matrices $\bm{Q}\in\mathbb{R}^{L_{hq}\times d_{qk}}$, $\bm{K}\in\mathbb{R}^{L_{kv}\times d_{qk}}$, $\bm{V}\in\mathbb{R}^{L_{kv}\times d_{hv}}$.
@@
2:Initialize $\bm{m}_{old},\,\bm{m}_{new}\in\mathbb{R}^{L_{q}}$ to -$\infty$ in SRAM.
@@
3:Initialize accumulators $\bm{\mathrm{H}}_{\text{intra}}\in\mathbb{R}^{B_{Lhq}\times B_{dv}}$ and $\bm{\mathrm{n}}^{(\text{intra})}\in\mathbb{R}^{B_{Lhq}}$ in SRAM.
4:Load $\bm{\mathrm{b}}_{k}^{(q)}\in\mathbb{R}^{B_{Lhq}}$ from HBM to SRAM.
5:for $i=1$ to $\left\lfloor\frac{(i_{Lq}+1)\cdot B_{Lhq}}{B_{Lkv}}\right\rfloor$ do
6:  Initialize accumulator $\bm{\mathrm{S}}^{(i)}\in\mathbb{R}^{B_{Lhq}\times B_{Lkv}}$ in SRAM.
7:  for $j=1$ to $\left\lceil\frac{d_{qk}}{B_{dqk}}\right\rceil$ do
8:   Load $\bm{Q}^{(j)}\in\mathbb{R}^{B_{Lhq}\times B_{dqk}}$ and $\bm{K}^{(j)}\in\mathbb{R}^{B_{Lkv}\times B_{dqk}}$ from HBM to SRAM.
9:   Accumulate $\bm{\mathrm{S}}^{(i)}\mathrel{+}=\bm{Q}^{(j)}{\bm{K}^{(j)}}^{\top}$.
10:  end for
11:  Load $\bm{\mathrm{b}}_{k}^{(kv)}\in\mathbb{R}^{B_{Lkv}}$ and $\bm{\mathrm{i}}_{k}^{(kv)}\in\mathbb{R}^{B_{Lkv}}$ from HBM to SRAM.
12:  Compute $\widetilde{\bm{\mathrm{D}}}^{(i)}=\bm{\mathrm{b}}_{k}^{(q)}-{\bm{\mathrm{b}}_{k}^{(kv)}}^{\top}+{\bm{\mathrm{i}}_{k}^{(kv)}}^{\top}\in\mathbb{R}^{B_{Lhq}\times B_{Lkv}}$.
13:  if $i\cdot B_{Lkv}\geqslant i_{Lq}\cdot B_{Lhq}$ then
14:   Apply causal mask to $\widetilde{\bm{\mathrm{D}}}^{(i)}$.
15:  end if
16:  Compute $\bm{m}_{new}=\mathrm{maximum}\{\bm{m}_{old},\,\mathrm{rowmax}\;\widetilde{\bm{\mathrm{D}}}^{(i)}\}$.
17:  Compute $\bm{\mathrm{D}}^{(i)}=\exp(\widetilde{\bm{\mathrm{D}}}^{(i)}-\bm{m}_{new})$.
18:  Compute $\overline{\bm{\mathrm{S}}}^{(i)}=\frac{1}{\sqrt{d_{qk}}}\bm{\mathrm{S}}\odot\bm{\mathrm{D}}^{(i)}$.
19:  Load $\bm{V}^{(i)}\in\mathbb{R}^{B_{Lkv}\times B_{dhv}}$ for Block $i_{dhv}$ from HBM to SRAM.
20:  Accumulate $\bm{\mathrm{H}}_{\text{intra}}=\exp(\bm{m}_{old}-\bm{m}_{new})\cdot\bm{\mathrm{H}}_{\text{intra}}+\overline{\bm{\mathrm{S}}}^{(i)}\,\bm{V}$.
21:  Accumulate $\bm{\mathrm{n}}^{(\text{intra})}=\exp(\bm{m}_{old}-\bm{m}_{new})\cdot\bm{\mathrm{n}}^{(\text{intra})}+\overline{\bm{\mathrm{S}}}^{(i)}\,\bm{1}$.
22:  Update $\bm{m}_{old}=\bm{m}_{new}$.
23:end for
@@
24:Load ${m}^{(\text{inter})}_{k-1}\in\mathbb{R}$ from HBM to SRAM.
25:Compute $\bm{m}^{(\text{combine})}_{k}=\mathrm{maximum}\left\{\bm{\mathrm{b}}_{k}^{(q)}+{m}^{(\text{inter})}_{k-1},\,\bm{m}_{new}\right\}$.
26:Compute $\overline{\bm{\mathrm{b}}}_{k}=\exp\left(\bm{\mathrm{b}}_{k}^{(q)}+{m}^{(\text{inter})}_{k-1}-\bm{m}^{(\text{combine})}_{k}\right)$.
27:Initialize accumulators $\bm{\mathrm{H}}_{\text{inter}}\in\mathbb{R}^{B_{Lhq}\times B_{dhv}}$ for Block $i_{dhv}$ and $\bm{\mathrm{n}}^{(\text{inter})}\in\mathbb{R}^{B_{Lhq}}$ in SRAM.
@@
28:for $j=1$ to $\left\lceil\frac{d_{qk}}{B_{dqk}}\right\rceil$ do
29:  Load $\bm{Q}^{(j)}\in\mathbb{R}^{B_{Lhq}\times B_{dqk}}$ and $\bm{C}_{k-1}^{(j)}\in\mathbb{R}^{B_{dqk}\times B_{dhv}}$ for Block $i_{dhv}$ from HBM to SRAM.
30:  Compute $\overline{\bm{Q}}^{(j)}=\frac{1}{\sqrt{d_{qk}}}\,\bm{Q}^{(j)}\odot\bm{\mathrm{b}}_{k}^{(q)}$.
31:  Accumulate $\bm{\mathrm{H}}_{\text{inter}}\mathrel{+}=\overline{\bm{Q}}^{(j)}\,\bm{C}_{k-1}^{(j)}$.
32:  Load $\bm{\bm{n}}_{k-1}^{(j)}\in\mathbb{R}^{B_{dqk}}$.
33:  Accumulate $\bm{\mathrm{n}}^{(\text{inter})}\mathrel{+}=\overline{\bm{Q}}^{(j)}\,\bm{\bm{n}}_{k-1}^{(j)}$.
[…]
35:Compute $\bm{\mathrm{H}}^{(\text{comb})}=\bm{\mathrm{H}}_{\text{intra}}+\exp(\bm{m}_{new}-\bm{m}^{(\text{combine})}_{k})\,\bm{\mathrm{H}}_{\text{inter}}$.
36:Compute $\bm{\mathrm{n}}^{(\text{comb})}=\bm{\mathrm{n}}^{(\text{intra})}+\exp(\bm{m}_{new}-\bm{m}^{(\text{combine})}_{k})\,\bm{\mathrm{n}}^{(\text{inter})}$.
37:Compute $\bm{\mathrm{H}}^{(k)}=\frac{\bm{\mathrm{H}}^{(\text{comb})}}{\max\left\{|\bm{\mathrm{n}}^{(\text{comb})}|,\exp(-\bm{m}^{(\text{combine})}_{k})\right\}}$.
38:Store $\bm{\mathrm{H}}^{(k)}$, $\bm{\mathrm{n}}^{(\text{comb})}$ and $\bm{m}^{(\text{combine})}_{k}$ to HBM.
C.4 TFLA Backward Pass
@@ D.1 Stabilization of the Exponential Input Gate
$\bm{h}_{t}=\tilde{\mathbf{o}}_{t}\odot\frac{\bm{C}_{t}^{\top}\,\bm{q}_{t}}{\max\left\{{{\left|\bm{n}_{t}^{\top}\ \bm{q}_{t}\right|}},\ 1\right\}},$ (96)
@@
$\widetilde{\bm{C}}_{t}$ $=\bm{C}_{t}\,\exp(-m_{t})\$$\rightarrow\ \widetilde{\bm{C}}_{t-1}=\bm{C}_{t-1}\,\exp(-m_{t-1})$ $\Leftrightarrow\ {\bm{C}}_{t-1}=\widetilde{\bm{C}}_{t-1}\,\exp(m_{t-1})$ (100)
$\widetilde{\bm{n}}_{t}$ $=\bm{n}_{t}\,\exp(-m_{t})\$$\rightarrow\ \widetilde{\bm{n}}_{t-1}\ =\bm{n}_{t-1}\,\exp(-m_{t-1})$ $\Leftrightarrow\ {\bm{n}}_{t-1}\ =\widetilde{\bm{n}}_{t-1}\,\exp(m_{t-1})$